# Estimation of gravitational production uncertainties

### v1

@@ Abstract
Cosmological production of scalar, non-minimally coupled dark matter depends on the specifics of the inflationary model under consideration. We analyze both Starobinsky inflation and a quadratic potential, solve the full background dynamics, study pair production during inflation and reheating, and find that the observed dark matter abundance can be explained solely by this mechanism, regardless of the inflationary model. Qualitative differences between the two cases only appear for dark matter masses close to the inflationary scale. In addition, we identify a large region in parameter space in which cosmological production of dark matter is mostly independent of the chosen inflationary potential, highlighting the robustness of this dark matter production mechanism and its independence of the unknown particular details of inflation. In the region of masses lower than the scale of inflation, and sufficiently away from the conformal limit, the total comoving number density of produced particles becomes a function of the coupling to the geometry alone. This allows us to provide an approximated analytic expression for fitting the resulting abundance.
IPARCOS-UCM-25-XXX
IFT-UAM/CSIC-25-XXX
@@ 1 Introduction
Notation. We set $\text{M}_{\text{P}}=1/\sqrt{G},\,\hbar=c=k_{B}=1$, and use the signature (,+,+,+)(-,+,+,+)( - , + , + , + ). We will use the quadratic inflation potential mass, fixed to $m_{\phi}=1.21\times 10^{13}\ \mathrm{GeV}$, as a reference energy scale for comparison between both models.
@@ 2.1 Dynamics of a scalar field in FLRW geometry
We will consider a massive scalar field $\varphi$ with non-minimal coupling to gravity through the Ricci curvature scalar $R$. For cosmological cases like the one at hand, the geometry is well described by the Friedmann-Lemaître-Robertson-Walker (FLRW) metric. For simplicity, and in agreement with observations [56], we consider flat spatial sections. The metric is therefore
${\text{d}}s^{2}=a^{2}(\eta)({\text{d}}\eta^{2}-{\text{d}}\bm{x}^{2}),$ (2.1)
in the so-called conformal time $\eta$, related to cosmological time $t$ by the relation $a(\eta){\text{d}}\eta={\text{d}}t$.
@@
The mode equation (2.4) has a two-dimensional space of complex solutions, and therefore the set {$v_{k},v^{*}_{k}\}$—for a given value of $k$—forms a basis of the space of solutions. An orthonormal basis compatible with the Klein-Gordon product associated with the equation of motion (2.2[17] fulfills the following condition on its Wronskian,
@@ 2.2 Quantization
$[a_{\bm{k}},a^{\dagger}_{\bm{k^{\prime}}}]=\delta^{(3)}(\bm{k}-\bm{k^{\prime}}),\qquad[a_{\bm{k}},a_{\bm{k^{\prime}}}]=0,\qquad[a^{\dagger}_{\bm{k}},a^{\dagger}_{\bm{k^{\prime}}}]=0.$ (2.9)
Thus, for a given basis of solutions $\{v_{k},v_{k}^{*}\}$, one can find creation and annihilation operators $a_{\bm{k}}$, $a_{\bm{k}}^{{\dagger}}$ and define a vacuum state |$\ket{0}$ such that $a_{\bm{k}}\ket{0}=0$ and particle states in the usual way. However, a different basis $\{u_{k},u_{k}^{*}\}$ will give another set of creation and annihilation operators $b_{\bm{k}}$, $b_{\bm{k}}^{\dagger}$—and thus to another definition for the vacuum and particle states. These new $b$-operators are related to the $a$-operators by the same Bogoliubov coefficients that relate the different set of modes, through
@@
One could, in principle, impose certain criteria for selecting one particular Fock quantization. For example, in Minkowski spacetime, Poincaré invariance alone selects a preferred quantum vacuum: the Minkowski vacuum. However, in general, spacetime symmetries are not enough select a unique quantization.
@@ 4.2 Choice of out vacuum
where the frequency $\omega_{k,\text{av}}$ is defined by averaging the Ricci scalar oscillations during reheating, using the fact that the Ricci amplitude decays as $1/t^{2}$ this regime. This allows us to write
@@
where $\langle R(\eta)\rangle=\mathcal{C}/t^{2}$ denotes the amplitude of the oscillations, and $\mathcal{C}$ a constant factor. More details on the validity of this prescription can be found in [20].
@@ 5 Spectra and abundance of produced particles
The total density of produced particles as defined in equation (2.12) is obtained after integrating the spectra. If we assume that the dark matter field does not interact with any other particles, and cosmological production becomes negligible from $\eta_{\text{f}}$ onward, the physical density today can be computed as $n/a_{\text{today}}^{3}$, taking into account the dilution due to the expansion of spacetime. In practice, we will compute the physical density at the end of reheating, $n/a_{\rm{rh}}^{3}$, and parametrize the dilution from $\eta_{\text{rh}}$ until today with the reheating temperature $T_{\text{rh}}$. The time at which reheating ends, $\eta_{\text{rh}}$, can be found for a given reheating temperature by taking into account that at this moment radiation dominates, and the Hubble rate can be written as
@@
Figure 4: Abundance in the present day of produced particles for each potential as a function of their mass $m$ and coupling $\xi$, for a reheating temperature of $T_{\text{rh}}=10^{13}\ \mathrm{GeV}$. The dashed line corresponds to the observed abundance of dark matter.
@@
The time $\eta_{\text{f}}$, at which the adiabatic regime is reached for each combination of parameters limits the earliest reheating time $\eta_{\text{rh}}$ we can consider, since we assume that the dynamics are the same until $\eta_{\text{f}}$. However, even if reheating ends before adiabaticity is reached, one can argue that the number density of particles produced after reheating is negligible, since spacetime expansion becomes even more adiabatic. Therefore, in practice, we only require $\eta_{\text{avg}}\leq\eta_{\text{rh}}$, which sets the highest reheating temperature allowed to $T_{\text{rh}}\sim 10^{13}\,\text{GeV}$.
@@
In figure 4 we show the abundances in the present day corresponding to the quadratic (left) and Starobinsky (right) models, as function of the dark matter mass $m$ and its coupling to the geometry $\xi$, for a reheating temperature of $T_{\text{rh}}=10^{13}\,\text{GeV}$. Crucially, the resulting abundance is very similar except in the $m\gtrsim m_{\phi},\bar{m}_{\phi}$ region, when mass of the spectator field is of the order of the inflaton mass. This is expected, as the typical scale of inflation is directly related to this quantity. When the mass of the dark matter candidate becomes of the order of the inflaton mass, production becomes inefficient. Since Starobinsky’s inflaton mass is larger, this behavior is expected at higher masses than in the quadratic case.

### v2

@@ Abstract
Gravitational production of scalar, non-minimally coupled dark matter depends on the specifics of the inflationary model under consideration. We analyze both Starobinsky inflation and a quadratic potential, solve the full background dynamics, study pair production during inflation and reheating, and find that the observed dark matter abundance can be explained solely by this mechanism, regardless of the inflationary model. Qualitative differences between the two cases only appear for dark matter masses close to the inflationary scale. In addition, we identify a large region in parameter space in which gravitational production of dark matter is mostly independent of the chosen inflationary potential, highlighting the robustness of this dark matter production mechanism and its independence of the unknown particular details of inflation. In the region of masses lower than the scale of inflation, and sufficiently away from the conformal limit, the total comoving number density of produced particles becomes a function of the coupling to the geometry alone. This allows us to provide an approximated analytic expression for fitting the resulting abundance.
IPARCOS-UCM-25-062
IFT-UAM/CSIC-25-73
@@ 1 Introduction
Notation. We set $\text{M}_{\text{P}}=1/\sqrt{G},\,\hbar=c=k_{B}=1$, and use the signature (,+,+,+)(-,+,+,+). We will use the quadratic inflation potential mass, fixed to $m_{\phi}=1.21\times 10^{13}\ \mathrm{GeV}$, as a reference energy scale for comparison between both models.
@@ 2.1 Dynamics of a scalar field in FLRW geometry
We will consider a massive scalar field $\varphi$ with non-minimal coupling to gravity through the Ricci curvature scalar $R$. Its corresponding action reads
$S=-\frac{1}{2}\int{\text{d}}^{4}x\sqrt{-g}(\partial_{\mu}\varphi\partial^{\mu}\varphi+m^{2}\varphi^{2}+\xi R\varphi^{2}),$ (2.1)
where $g$ is the determinant of the metric, $m$ is the field mass, and $\xi$ denotes the strength of the curvature coupling. For cosmological cases like the one at hand, the geometry is well described by the Friedmann-Lemaître-Robertson-Walker (FLRW) metric. For simplicity, and in agreement with observations [56], we consider flat spatial sections. The metric is therefore
${\text{d}}s^{2}=a^{2}(\eta)(-{\text{d}}\eta^{2}+{\text{d}}\bm{x}^{2}),$ (2.2)
in the so-called conformal time $\eta$, related to cosmological time $t$ by the relation $a(\eta){\text{d}}\eta\equiv{\text{d}}t$.
@@
The mode equation (2.5) has a two-dimensional space of complex solutions, and therefore the set $v_{k},v^{*}_{k}\}$—for a given value of kk—forms a basis of the space of solutions. An orthonormal basis compatible with the Klein-Gordon product associated with the equation of motion (2.3[17] fulfills the following condition on its Wronskian,
@@ 2.2 Quantization
$a_{\bm{k}},a^{\dagger}_{\bm{k^{\prime}}}]=\delta^{(3)}(\bm{k}-\bm{k^{\prime}}),\qquad[a_{\bm{k}},a_{\bm{k^{\prime}}}]=0,\qquad[a^{\dagger}_{\bm{k}},a^{\dagger}_{\bm{k^{\prime}}}]=0.$ (2.10)
[…]
@@
One could, in principle, impose certain criteria for selecting one particular Fock quantization. For example, in Minkowski spacetime, Poincaré invariance alone selects a preferred quantum vacuum: the Minkowski vacuum. However, in general, spacetime symmetries are not enough to select a unique quantization.
@@ 4.2 Choice of out vacuum
where the frequency $\omega_{k,\text{av}}$ is defined by averaging the Ricci scalar oscillations during reheating, using the fact that the Ricci amplitude decays as $1/t^{2}$ within this regime. This allows us to write
@@
where $\langle R(\eta)\rangle=\mathcal{C}/t^{2}(\eta)$ denotes the amplitude of the oscillations well inside the reheating epoch, and $\mathcal{C}$ a constant factor. For each potential, $\mathcal{C}$ is calculated by evaluating the Ricci scalar at a late-time $\eta_{\mathcal{C}}$ corresponding to a maximum of the oscillation, so that $\langle R(\eta_{\mathcal{C}})\rangle\equiv R(\eta_{\mathcal{C}})$.
In practice, we solve the mode equation equation (2.5) until a time $\bar{\eta}$ for all pairs $(m,\xi)$ in the considered parameter space. If $\lvert\omega_{k}^{\prime}/\omega_{k}^{2}\rvert\leq 0.01$ at that time for all the wavenumbers $k$ of interest (and therefore condition (4.4) is fulfilled), we use the adiabatic prescription, identifying $\bar{\eta}$ with $\eta_{\text{f}}$. Otherwise, we use the averaged vacuum prescription and identify $\bar{\eta}$ with $\eta_{\text{av}}$. The values of $\bar{\eta}$ are $1.6\times 10^{3}m_{\phi}$ for the quadratic scenario and $3.6\times 10^{2}m_{\phi}$ for the Starobinsky potential, leading to $\mathcal{C}=5.4$ and $\mathcal{C}=12.8$, respectively. To ensure that the averaged vacuum provides the right number density of produced particles, we consider the worst case scenario—i.e., the values of $m$ and $\xi$ leading to the least adiabatic frequency—and compare both results. Let us stress that this is nothing but a method for obtaining the number density of produced particles at $\eta_{\text{f}}$, which is the physical quantity we are interested in. More details on the validity of this prescription can be found in [20].
@@ 5 Spectra and abundance of produced particles
The total density of produced particles as defined in equation (2.13) is obtained after integrating the spectra. If we assume that the dark matter field does not interact with any other particles, and gravitational production becomes negligible from $\eta_{\text{f}}$ onward, the physical density today can be computed as $n/a_{\text{today}}^{3}$, taking into account the dilution due to the expansion of spacetime. In practice, we will compute the physical density at the end of reheating, $n/a_{\rm{rh}}^{3}$, and parametrize the dilution from $\eta_{\text{rh}}$ until today with the reheating temperature $T_{\text{rh}}$. The time at which reheating ends, $\eta_{\text{rh}}$, can be found for a given reheating temperature by taking into account that at this moment radiation dominates, and the Hubble rate can be written as
@@
Figure 4: Abundance in the present day of produced particles for each potential as a function of their mass $m$ and coupling $\xi$, for a reheating temperature of $T_{\text{rh}}=10^{13}\ \mathrm{GeV}$. The dashed line corresponds to the observed abundance of dark.
@@
The time $\eta_{\text{f}}$, at which the adiabatic regime is reached for each combination of parameters limits the earliest reheating time $\eta_{\text{rh}}$ we can consider, since we assume that the dynamics are the same until $\eta_{\text{f}}$. However, even if reheating ends before adiabaticity is reached, one can argue that the number density of particles produced after reheating is negligible, since spacetime expansion becomes even more adiabatic. Therefore, in practice, we only require $\eta_{\text{}}\leq\eta_{\text{rh}}$, which sets the highest reheating temperature allowed to $T_{\text{rh}}\sim 10^{13}\,\text{GeV}$.
@@
In Figure 4 we show the abundances in the present day corresponding to the quadratic (left) and Starobinsky (right) models, as function of the dark matter mass $m$ and its coupling to the geometry $\xi$, for a reheating temperature of $T_{\text{rh}}=10^{13}\,\text{GeV}$. Note that we consider abundances larger than the observed amount of dark matter, as other non-standard cosmological evolutions could further dilute the number density of produced particles. Then, the observed amount of dark matter could be recovered in this region of parameter space. Crucially, the resulting abundance is very similar except in the $m\gtrsim m_{\phi},\bar{m}_{\phi}$ region, when mass of the spectator field is of the order of the inflaton mass. This is expected, as the typical scale of inflation is directly related to this quantity. When the mass of the dark matter candidate becomes of the order of the inflaton mass, production becomes inefficient. Since Starobinsky’s inflaton mass is larger, this behavior is expected at higher masses than in the quadratic case.
@@ 6 Conclusions
In the future, we aim to investigate other types of spectator fields to test how particle production depends on the underlying inflationary dynamics. While some modes exhibit more complex behavior—such as the longitudinal mode of a vector field [44]—, others, like the transverse modes of vector fields, behave analogously to scalar fields. In such cases, we expect our conclusions to carry over without significant modification. A more comprehensive analysis would require to also examine qualitatively different inflationary potentials, especially those that do not behave quadratically during reheating. Although definitive statements require explicit calculations, we expect that for conformally coupled fields, pair production proceeds in a similar fashion, as curvature oscillations play no significant role in this case.